OpenStudy (anonymous):

How do I solve the system? x^2 + xy = 1 xy + y^2 = 3

OpenStudy (danjs):

you can maybe solve one of them for either x or y, and put all that in for the variable in the other equation, to reduce to single var

jimthompson5910 (jim_thompson5910):

Solve for `xy` in either equation. I'm going to pick on the second equation xy + y^2 = 3 xy + y^2-y^2 = 3-y^2 xy = 3-y^2 then plug this into the other equation x^2 + xy = 1 x^2 + 3-y^2 = 1 x^2-y^2 + 3 = 1 x^2-y^2 + 3-3 = 1-3 x^2 - y^2 = -2 This is as far as you can go really. We'd need another equation to figure out what x and y are. In this case, there are infinitely many solutions. Each solution (x,y) lies along a hyperola

OpenStudy (danjs):

plug in (1/2 , 3/2)

OpenStudy (arnavguddu):

x.x+xy=x(x+y)=1 xy+y.y=y(x+y)=3 divide first eqn. by second, x/y=1/3 or y=3x now solve 1st eqn by substitution and solve for y, then find x

jimthompson5910 (jim_thompson5910):

@arnavguddu has the right idea. There will be 2 solutions in the form (x,y)

OpenStudy (danjs):

yeah that is good, y = 1/x - x solving the first equation for the y, and using that, you can get to the end too

OpenStudy (danjs):

the process you did @jim_thompson5910 looks like the same as adding (-1)times the first equation to the other eq does that linear system work for those nonlinear things too

OpenStudy (danjs):

linear sytem solving rules

OpenStudy (danjs):

constant multioplication and adding equations and changing order, linear algebra